# Compressed sensing quantum state tomography for qudits: A comparison of Gell-Mann and Heisenberg-Weyl observable bases

### v1

@@ III.3.2 Fixed Hilbert space dimension
Finally, we examine how the choice of quantum information unit—qubit or qudit—affects the efficiency of CS-QST when the Hilbert space dimension $d$ is fixed. By varying $N$ and $k$ while keeping $d=k^{N}$ constant, we compare the relative advantages of different representations. Specifically, we focus on the case of a density matrix with $d=16$, and analyze the following three configurations:: (i) a four-qubit system using the Pauli basis; (ii) a two-qudit system with $k=4$, using either the GGM or HWO basis; and (iii) a single-qudit system with $k=16$, using either the GGM or HWO basis.
@@ V.1 GGM and HW-observable operators for $k=4$
$\begin{split}&W(0,0)=\begin{pmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&1&0\\ 0&0&0&1\end{pmatrix}\hskip 5.69054ptW(0,1)=\begin{pmatrix}0&\chi^{*}&0&\chi\\ \chi&0&\chi^{*}&0\\ 0&\chi&0&\chi^{*}\\ \chi^{*}&0&\chi&0\end{pmatrix}\\ &W(0,2)=\begin{pmatrix}0&0&1&0\\ 0&0&0&1\\ 1&0&0&0\\ 0&1&0&0\end{pmatrix}\hskip 5.69054ptW(0,3)=\begin{pmatrix}0&\chi&0&\chi^{*}\\ \chi^{*}&0&\chi&0\\ 0&\chi^{*}&0&\chi\\ \chi&0&\chi^{*}&0\end{pmatrix}\\ &W(1,0)=\begin{pmatrix}1&0&0&0\\ 0&-1&0&0\\ 0&0&-1&0\\ 0&0&0&1\end{pmatrix}\hskip 5.69054ptW(1,1)=\frac{1}{\sqrt{2}}\begin{pmatrix}0&-i&0&1\\ i&0&-1&0\\ 0&-1&0&i\\ 1&0&-i&0\end{pmatrix}\\ &W(1,2)=\begin{pmatrix}0&0&-i&0\\ 0&0&0&i\\ i&0&0&0\\ 0&-i&0&0\end{pmatrix}\hskip 5.69054ptW(1,3)=\frac{1}{\sqrt{2}}\begin{pmatrix}0&-i&0&-1\\ i&0&1&0\\ 0&1&0&i\\ -1&0&-i&0\end{pmatrix}\\ &W(2,0)=\begin{pmatrix}1&0&0&0\\ 0&-1&0&0\\ 0&0&1&0\\ 0&0&0&-1\end{pmatrix}\hskip 5.69054ptW(2,1)=\begin{pmatrix}0&-\chi&0&\chi^{*}\\ -\chi^{*}&0&\chi&0\\ 0&\chi^{*}&0&-\chi\\ \chi&0&-\chi^{*}&0\end{pmatrix}\\ &W(2,2)=\begin{pmatrix}0&0&-1&0\\ 0&0&0&1\\ -1&0&0&0\\ 0&1&0&0\end{pmatrix}\hskip 5.69054ptW(2,3)=\begin{pmatrix}0&-\chi^{*}&0&\chi\\ -\chi&0&\chi^{*}&0\\ 0&\chi&0&-\chi^{*}\\ \chi^{*}&0&-\chi&0\end{pmatrix}\\ &W(3,0)=\begin{pmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&-1&0\\ 0&0&0&-1\end{pmatrix}\hskip 5.69054ptW(3,1)=\frac{1}{\sqrt{2}}\begin{pmatrix}0&-1&0&-i\\ -1&0&-i&0\\ 0&i&0&1\\ i&0&1&0\end{pmatrix}\\ &W(3,2)=\begin{pmatrix}0&0&i&0\\ 0&0&0&i\\ -i&0&0&0\\ 0&-i&0&0\end{pmatrix}\hskip 5.69054ptW(3,3)=\frac{1}{\sqrt{2}}\begin{pmatrix}0&1&0&-i\\ 1&0&-i&0\\ 0&i&0&-1\\ i&0&-1&0\end{pmatrix}.\end{split}$
@@ V.4 Model selection using AIC
The results of applying the AIC to various fitting curves using the methodology described above are presented in Table 2. Among the candidate models, the one with the lowest AIC value was selected as the optimal fitting model. The corresponding fitting curves are shown in Fig. 3(a), while the same curves divided by $d^{2}$ are presented in Fig. 3(b).

### v2

@@ III.3.2 Fixed Hilbert space dimension
Finally, we examine how the choice of quantum information unit—qubit or qudit—affects the efficiency of CS-QST when the Hilbert space dimension $d$ is fixed. By varying $N$ and $k$ while keeping $d=k^{N}$ constant, we compare the relative advantages of different representations. Specifically, we focus on the case of a density matrix with $d=16$, and analyze the following three configurations: (i) a four-qubit system using the Pauli basis; (ii) a two-qudit system with $k=4$, using either the GGM or HWO basis; and (iii) a single-qudit system with $k=16$, using either the GGM or HWO basis.
@@ V.1 GGM and HW-observable operators for $k=4$
$\begin{split}&W(0,0)=\begin{pmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&1&0\\ 0&0&0&1\end{pmatrix},\quad W(0,1)=\begin{pmatrix}0&\chi^{*}&0&\chi\\ \chi&0&\chi^{*}&0\\ 0&\chi&0&\chi^{*}\\ \chi^{*}&0&\chi&0\end{pmatrix},\quad W(0,2)=\begin{pmatrix}0&0&1&0\\ 0&0&0&1\\ 1&0&0&0\\ 0&1&0&0\end{pmatrix}\\ &W(0,3)=\begin{pmatrix}0&\chi&0&\chi^{*}\\ \chi^{*}&0&\chi&0\\ 0&\chi^{*}&0&\chi\\ \chi&0&\chi^{*}&0\end{pmatrix},\quad W(1,0)=\begin{pmatrix}1&0&0&0\\ 0&-1&0&0\\ 0&0&-1&0\\ 0&0&0&1\end{pmatrix},\quad W(1,1)=\frac{1}{\sqrt{2}}\begin{pmatrix}0&-i&0&1\\ i&0&-1&0\\ 0&-1&0&i\\ 1&0&-i&0\end{pmatrix},\\ &W(1,2)=\begin{pmatrix}0&0&-i&0\\ 0&0&0&i\\ i&0&0&0\\ 0&-i&0&0\end{pmatrix},\quad W(1,3)=\frac{1}{\sqrt{2}}\begin{pmatrix}0&-i&0&-1\\ i&0&1&0\\ 0&1&0&i\\ -1&0&-i&0\end{pmatrix},\quad W(2,0)=\begin{pmatrix}1&0&0&0\\ 0&-1&0&0\\ 0&0&1&0\\ 0&0&0&-1\end{pmatrix},\\ &W(2,1)=\begin{pmatrix}0&-\chi&0&\chi^{*}\\ -\chi^{*}&0&\chi&0\\ 0&\chi^{*}&0&-\chi\\ \chi&0&-\chi^{*}&0\end{pmatrix},\quad W(2,2)=\begin{pmatrix}0&0&-1&0\\ 0&0&0&1\\ -1&0&0&0\\ 0&1&0&0\end{pmatrix},\quad W(2,3)=\begin{pmatrix}0&-\chi^{*}&0&\chi\\ -\chi&0&\chi^{*}&0\\ 0&\chi&0&-\chi^{*}\\ \chi^{*}&0&-\chi&0\end{pmatrix},\\ &W(3,0)=\begin{pmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&-1&0\\ 0&0&0&-1\end{pmatrix},\quad W(3,1)=\frac{1}{\sqrt{2}}\begin{pmatrix}0&-1&0&-i\\ -1&0&-i&0\\ 0&i&0&1\\ i&0&1&0\end{pmatrix},\quad W(3,2)=\begin{pmatrix}0&0&i&0\\ 0&0&0&i\\ -i&0&0&0\\ 0&-i&0&0\end{pmatrix},\\ &W(3,3)=\frac{1}{\sqrt{2}}\begin{pmatrix}0&1&0&-i\\ 1&0&-i&0\\ 0&i&0&-1\\ i&0&-1&0\end{pmatrix}.\end{split}$
@@ V.4 Model selection using AIC
The results of applying the AIC to various fitting curves using the methodology described above are presented in Table 2. Among the candidate models, the one with the lowest AIC value was selected as the optimal fitting model. The corresponding fitting curves are shown in Fig. 3(a), while the same curves multiplied by $d^{2}$ are presented in Fig. 3(b).